If each side of a square is increased by 3 cm, the total area of both the new square and the original square will be 617 cm2 . Approximate to the nearest tenth of a centimeter the length of each side of the original square.
Answer by mananth(16946)

let side of square be x
area = l*w= x^2
3 is added to its side
Area of this square = (x+3)^2
x^2+(x+3)^2= 617
x^2+x^2+6x+9=617
2x^2+6x-617+9=0
2x^2+6x-608=0
/2
x^2+3x-304 =0
x^2+19x-16x-304=0
x(x+19)-16(x+19)=0
(x+19)(x-16)=0
x=-19 OR 16
ignore negative
the length of side = 16cm